# Enhancing Multimodal Medical Image Classification using Cross-Graph Modal Contrastive Learning

### v2

@@ Abstract
The classification of medical images is a pivotal aspect of disease diagnosis, often enhanced by deep learning techniques. However, traditional approaches typically focus on unimodal medical image data, neglecting the integration of diverse non-image patient data. This paper proposes a novel Cross-Graph Modal Contrastive Learning (CGMCL) framework for multimodal medical image classification. The model effectively integrates both image and non-image data by constructing cross-modality graphs and leveraging contrastive learning to align multimodal features in a shared latent space. An inter-modality feature scaling module further optimizes the representation learning process by reducing the gap between heterogeneous modalities. The proposed approach is evaluated on two datasets: a Parkinson’s disease (PD) dataset and a public melanoma dataset. Results demonstrate that CGMCL outperforms conventional unimodal methods in accuracy, interpretability, and early disease prediction. Additionally, the method shows superior performance in multi-class melanoma classification. The CGMCL framework provides valuable insights into medical image classification while offering improved disease interpretability and predictive capabilities.
@@
1 Introduction
In recent years, medical computing has increasingly shifted toward utilizing deep learning frameworks as a foundation for diagnostic analysis. Single convolutional neural networks (CNNs) have demonstrated substantial success in various unimodal medical imaging applications, such as computer-aided detection and diagnosis, image segmentation, and survival analysis [1, 2, 3]. However, the landscape of medical diagnostics has become increasingly complex with the growing availability of diverse data modalities. These include unstructured electronic health records (EHRs), quantitative blood test results, and demographic data, making it difficult to rely solely on single-modality data (e.g., clinical notes) for accurate disease predictions. [4, 5].
Injecting radioactive tracers for single-photon emission computed tomography (SPECT) scans remains a critical diagnostic tool for neurodegenerative diseases like Parkinson’s disease (PD), as it provides insights into the physiological functions and metabolic activities of organs beyond their anatomical structures [6]. However, SPECT has limitations, including relatively low spatial resolution and higher image noise due to the limited dose of radioactive tracers. Recent research has employed CNN-based models to address these issues, showing high accuracy in classifying and predicting PD [7, 8, 6]. Despite their success, these models often fail to incorporate or provide insights into relevant patient demographics in clinical practice.
Graph-based approaches offer a promising avenue for identifying commonalities among patients’ medical imaging data and symptom profiles [9, 10, 11]. Specifically, graph convolutional neural networks (GCNs) [12] have demonstrated remarkable efficacy in processing the complex interconnections and structural patterns inherent in non-linear data structures like graphs [11, 13]. While traditional machine learning methods often struggle with such sophisticated data, GCNs excel at interpreting and analyzing these relationships. GCNs learn from unstructured features by processing non-Euclidean distances, enabling them to perform various downstream tasks, including node classification, link prediction, and graph classification. An advanced framework of GCNs includes graph attention networks (GATs) [14], which introduce an attention mechanism to enhance the learning and fusion of neighboring node features. GATs can capture correlations between modalities, establish adaptive graph learning from heterogeneous multimodal feature spaces, and ultimately perform disease prediction in downstream tasks [15].
@@ 1 Introduction
Contrastive learning, a self-supervised framework, aims to minimize the distance between representations of positive sample pairs while maximizing it for negative pairs. This approach offers a promising solution to handling multimodal data similarity, complicated by the heterogeneous structural characteristics across different modalities [16, 17]. This approach aligns multimodal data in a shared embedding space for Alzheimer’s disease (AD) prediction. Additionally, attention mechanisms applied to tabular data improve model performance and interpretability [18], while multimodal transformers that combine image and clinical data can predict AD progression. [19]. However, these studies have typically focused on a specific disease, with limited research on model generalization across different multimodal disease datasets. This study proposes a novel approach that leverages dual graph networks to construct a cross-graph modal feature topology for diverse modalities, such as medical images and quantitative parameters. Our method employs GATs to learn feature encodings across modalities through node embedding. Additionally, we design a graph contrastive loss to enhance the similarity of the final representations. We validate our proposed Cross-Graph Modal Contrastive Learning (CGMCL) approach for multimodal fusion as shown in Fig. 2, and clinical interpretability using two multimodal medical datasets: a private PD dataset and a public melanoma dataset.
@@
2 RELATED WORK
2.1 Multimodal Learning in Medical Imaging
Image-based models have traditionally relied on unimodal input, primarily for disease classification. However, structured medical images alone are often insufficient for integrating patients’ physiological or numerical characteristics. Recent research has demonstrated the effectiveness of multimodal deep-learning approaches. For instance, a multi-co-attention model successfully integrated brain SPECT images with DNA methylation data [20]. In multimodal survival analysis, hazard functions are estimated by integrating features from co-attention modules applied to diverse pathology images and genetic data [2, 21]. However, in the above studies, feature extraction from medical images has predominantly been conducted at the pixel level, with limited consideration for structured feature learning that accounts for inter-patient image similarities (e.g., graph structures). Therefore, we propose that early diagnosis, particularly for conditions such as PD, requires more precise multimodal models to improve early prediction and clinical interpretability.
2.2 Multi-Modality Fusion Methods
The greatest challenge lies in effectively fusing modalities from different domains. Fig. 1 illustrates various feature fusion methods, including feature extraction from two modalities, followed by fusion approaches such as vector concatenation, attention-based fusion, and joint network fusion learning to combine cross-domain features [22], as shown in Fig. 1 (a)-(c). However, those fusion methods do not account for calculating similarities between different modalities. We first construct a graph structure using image feature vectors extracted from a pre-trained ResNet model as single-modality input features for each graph node to address this. For example, a GCN can use patch images from areas of pathological interest after gridding, with nodes representing patients and edges representing correlations between them [23, 24]. In this study, we propose a cross-graph model that inputs features from two separate encoder modalities. Subsequently, we construct a common modality embedding to better fuse the two domain latent spaces, as illustrated in Fig. 1 (d).
3 METHODOLOGY
3.1 Problem Definition
This study focuses on two medical modalities: medical images and non-image data (e.g., meta-features). We denote a multimodal set $\textbf{X}=\{X_{1},X_{2},...,X_{N}\}$ for $N$ patients. We can represent the data for the $i$-th patient as a 3-tuple $X_{i}=\{(I_{i},C_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}\in\mathbb{R}^{h\times h}$ represents the medical image with a scale dimension $h$, $C_{i}\in\mathbb{R}^{F}$ represents the meta-features with $F$ features, and $Y_{i}$ corresponds to the disease labels. We first construct a non-linear model (e.g., an CNN) to generate initial feature maps $I_{i}^{\prime}$ from the images:
@@ 3.1 Problem Definition
3.2 Graph Construction
In practice, to better measure the structural properties of the two modalities in non-Euclidean distance feature spaces, we consider the two modality graphs $\mathcal{G}^{I}(\mathcal{E}^{I},\mathcal{V}^{I})$ and $\mathcal{G}^{C}(\mathcal{E}^{C},\mathcal{V}^{C})$, with edges $\left|\mathcal{E}^{I}\right|$ and $\left|\mathcal{E}^{C}\right|$ and vertices $\left|\mathcal{V}^{I}\right|$ and $\left|\mathcal{V}^{C}\right|$, respectively. Given the input encoded features $F_{i}$ (e.g., $I_{i}^{\prime}$ or $C_{i}$), we can construct the binary adjacency matrices $A^{I}$ and $A^{C}$ using a $K$-nearest neighbors graph [25]:
@@ 3.2 Graph Construction
3.3 Graph Attention Encoder
Our modality-based graph model employs a GAT as the foundational features encoder to facilitate cross-graph modal learning across different graph modalities. Given the two input features node representations $I^{\prime}=[I^{\prime}_{1},I^{\prime}_{2},...,I^{\prime}_{N}]\in\mathbb{R}^{N\times F^{\prime}}$ and $C=[C_{1},C_{2},...,C_{N}]\in\mathbb{R}^{N\times F^{\prime}}$, we can formulate the multi-modality attention coefficients in a GNN as:
@@ 3.3 Graph Attention Encoder
3.4 Inter-Modality Feature Enhancement and Scaling
@@
3.5 Cross-Graph Modal Fusion
To integrate features from cross-graph modality representations at varying scales, we concatenate each image embedding $E^{I}$ with its corresponding representation $\mathcal{H}^{I}$ to form $\hat{Z}^{I}=[\mathcal{H}^{I}\mathbin{\|}E^{I}]$, and similarly for clinical data, $\hat{Z}^{C}=[\mathcal{H}^{C}\mathbin{\|}E^{C}]$, resulting in final embedding matrices. We then combine these cross-modal embeddings by element-wise addition and apply a sigmoid function: $\hat{Z}=\sigma(\hat{Z}^{I}+\hat{Z}^{C})$. This process aligns representations from both modalities in a common feature space. Furthermore, to integrate the feature spaces of images and meta-features in the same embedded space, we construct a similarity matrix $\textbf{S}\in\mathbb{R}^{N\times N}$ for each pair of similar patients using the final embedding $\hat{Z}$ learned from the model. The similarity between the $i$-th and $j$-th patients can be defined as follows:
@@ 3.5 Cross-Graph Modal Fusion
3.6 Contrastive Loss Optimization
@@ 4 Dataset Collection
5 EXPERIMENTS
5.1 Baseline Methods Comparison
@@
5.2 Compare with the SOTA Methods
@@ 6 Ablation Study
6.1 Parameters of $\beta$ in the Objective Function
@@
6.2 Multimodal Features Alignment in KL Divergence Converge
@@ 6.4 Objective Function Ablation
7 Visualization of Multimodal Representation
7.1 Identification of PD subtype trajectories
7.2 Multi-Class Trajectories in Melanoma Embedding
@@
8 Discussion
8.1 Clustering Results
@@
8.2 Neurodegenerative Subtype Analysis in PD
@@
8.3 Explainable Modality Diagnosis

### v4

@@ Abstract
The classification of medical images is a pivotal aspect of disease diagnosis, often enhanced by deep learning techniques. However, traditional approaches typically focus on unimodal medical image data, neglecting the integration of diverse non-image patient data. This paper proposes a novel Cross-Graph Modal Contrastive Learning (CGMCL) framework for multimodal structured data from different data domains to improve medical image classification. The model effectively integrates both image and non-image data by constructing cross-modality graphs and leveraging contrastive learning to align multimodal features in a shared latent space. An inter-modality feature scaling module further optimizes the representation learning process by reducing the gap between heterogeneous modalities. The proposed approach is evaluated on two datasets: a Parkinson’s disease (PD) dataset and a public melanoma dataset. Results demonstrate that CGMCL outperforms conventional unimodal methods in accuracy, interpretability, and early disease prediction. Additionally, the method shows superior performance in multi-class melanoma classification. The CGMCL framework provides valuable insights into medical image classification while offering improved disease interpretability and predictive capabilities. The code and data are available from https://github.com/Ding1119/CGMCL
@@
I Introduction
In recent years, medical computing has increasingly shifted toward utilizing deep learning frameworks as a foundation for diagnostic analysis. In particular, multimodal deep learning has emerged as a powerful approach for integrating diverse data sources to enhance predictive accuracy in various medical applications. The complex and heterogeneous nature of medical data, including radiological images, clinical records and histological images, necessitates the use of multimodal fusion methods to extract complementary features and improve clinical decision-making [1, 2, 3]. However, in real-world clinical settings, many modalities show inconsistency in modality distributions depending on disease characteristics, which indirectly creates limitations for multimodal learning [4]. Addressing these challenges, some studies have proposed advanced fusion architectures, such as feature-level fusion, decision-level fusion, and hybrid fusion, to optimize the integration of multimodal information [2, 5]. For example, integrating dermatological images with patient metadata significantly enhances the classification accuracy of skin cancer detection systems, outperforming unimodal image-based models [5].
Graph-based deep learning such as graph convolutional neural networks (GCNs) [6] learn from unstructured features by processing non-Euclidean distances, enabling them to perform various downstream tasks, including node classification, link prediction, and graph classification. An advanced framework of GCNs offer a promising avenue for identifying commonalities among patients in disease relationships via medical imaging data or symptom similarity [7, 8]. However, these methods only perform disease prediction on a single modality and lack information from both images and meta-features. AMA-GCN [9] uses multi-layer aggregation mechanisms and automatic feature selection to integrate both image and non-image data for various diseases. Nevertheless, AMA-GCN’s thresholds and aggregation methods require readjustment, and the model may overly depend on certain features, potentially leading to overfitting. Moreover, there is no unified approach for effectively achieving clinical interpretability through multimodal fusion, particularly for images and metadata associated with distinct graph structures in clinical settings, which should be further leveraged to enhance efficient latent space learning [10].
Recent studies have focused on aligning multimodal data in a shared embedding space for Alzheimer’s disease (AD) prediction. Notably, attention mechanisms applied to tabular data have demonstrated improved model performance and interpretability [11], while multimodal transformers combining image and clinical data have shown success in predicting AD progression [12]. Furthermore,  [13] has demonstrated the effectiveness of self-supervised contrastive learning in handling different modalities through dedicated encoder networks, which map data into a shared feature space while optimizing a contrastive loss function. However, these studies have primarily focused on specific diseases, with limited exploration of model generalization across diverse multimodal disease datasets and a lack of meta-features interpretable for disease classification tasks.
@@ I Introduction
To address above problems, we proposes a novel approach that leverages dual graph networks to construct a cross-graph modal feature topology for diverse modalities, such as medical images and quantitative parameters. Our method employs graph attention networks (GATs) [14] to learn feature encodings across modalities through node embedding. Additionally, we design a graph contrastive loss to enhance the similarity of the final representations. We validate our proposed Cross-Graph Modal Contrastive Learning (CGMCL) approach for multimodal fusion, and clinical interpretability using two multimodal medical datasets: a private PD dataset and a public melanoma dataset. The main contributions of this study can be summarized as follows:
  • We propose a multimodal fusion framework incorporating multi-structured modalities (e.g., medical images and metadata) from different patient disease characteristics as graph-structured features to enhance classification accuracy through scale and eliminate multimodality distribution inconsistent gaps.
  • To better evaluate the robustness of our proposed CGMCL model, we achieved more effective performance compared to fashion CNN-based models across two different types of multimodal datasets.
  • By incorporating meta-features into CGMCL, we enhance the model’s interpretability for neurodegenerative disease diagnosis.
II RELATED WORK
II-A Multimodal Learning in in Medical Imaging
Traditional image-based models approaches often fail to capture the full complexity of diseases, whereas multimodal fusion leverages complementary information to enhance predictive performance [15]. Some non-deep learning research works focus on using Wavelet Transform for multimodal image fusion in MRI/CT, PET/CT, and PET/MRI [16, 17, 18]. In neurodegenerative diseases diagnosis, studies have demonstrated that combining T1-weighted MRI, PET, and cerebrospinal fluid (CSF) biomarkers significantly improves early detection and progression monitoring of Parkinson’s Disease (PD) and Alzheimer’s Disease (AD) [19, 15]. Moreover, recent multimodal deep learning research focuses on PD classification applied multi-co-attention model successfully integrated brain single-photon emission computed tomography (SPECT) images with DNA methylation data [20]. However, in the above studies, feature extraction from medical images has predominantly been conducted at the pixel level, with limited consideration for structured feature learning that accounts for inter-patient image similarities (e.g., graph structures). Therefore, we propose that early diagnosis, particularly for conditions such as PD, requires more precise multimodal models to improve early prediction and clinical interpretability.
II-B Multi-Modality Fusion Strategies
According to deep learning fusion strategies, including feature extraction from two modalities, followed by fusion approaches such as vector concatenation, attention-based fusion, and joint network fusion learning to combine cross-domain features [21, 22, 23, 24, 21], as shown in Fig. 1 (a)-(c). Recent advancements in complicated multimodal framework have introduced more sophisticated fusion techniques, such as transformer-based fusion. MATR has demonstrated superior performance in multimodal medical image fusion by leveraging adaptive convolution and transformer mechanisms to utilize local and global contextual information [25]. However, the greatest challenge with the approach mentioned above is that it does not consider the different characteristics of different modalities when performing modality fusion, which creates gaps in feature representation between their distributions.
To learn consistent latent representations, we designed a dual-graph architecture. The first graph utilizes patient image features extracted from a pre-trained ResNet model as node representations, where each node represents a patient. The second graph incorporates structured meta-features, as illustrated in Fig. 1 (d). We then employ a dual-structured graph attention-based fusion framework to project and scale these different modalities into a consistent feature space, as shown in Fig. 2
@@ II-B Multi-Modality Fusion Strategies
III METHODOLOGY
III-A Problem Definition and Formulation
In this section, we formulate similarities within a single modality across different patients as graph structures to enhance disease classification compared to unimodal approaches. This study focuses on two medical modalities: medical images and non-image data (e.g., meta-features). We denote a multimodal set $\textbf{X}=\{X_{1},X_{2},...,X_{N}\}$ for $N$ patients. We can represent the data for the $i$-th patient as a 3-tuple $X_{i}=\{(I_{i},C_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}\in\mathbb{R}^{h\times h}$ represents the medical image with a scale dimension $h$, $C_{i}\in\mathbb{R}^{F}$ represents the meta-features with $F$ features, and $Y_{i}$ corresponds to the disease labels. We first construct a non-linear model (e.g., an CNN) to generate initial feature maps $I_{i}^{\prime}$ from the images:
@@ III-A Problem Definition and Formulation
III-B Graph Construction
In practice, to better measure the structural properties of the two modalities for each patient in non-Euclidean distance feature spaces, we consider the two modality graphs, structured-image graph $\mathcal{G}^{I}(\mathcal{E}^{I},\mathcal{V}^{I})$ and structured meta-feature graph $\mathcal{G}^{C}(\mathcal{E}^{C},\mathcal{V}^{C})$, with edges $\left|\mathcal{E}^{I}\right|$ and $\left|\mathcal{E}^{C}\right|$ and vertices $\left|\mathcal{V}^{I}\right|$ and $\left|\mathcal{V}^{C}\right|$, respectively. Given the input encoded features $F_{i}$ (e.g., $I_{i}^{\prime}$ or $C_{i}$), we can construct the binary adjacency matrices $A^{I}$ and $A^{C}$ using a $K$-nearest neighbors graph [26]:
@@ III-B Graph Construction
III-C Graph Attention Encoder
Our modality-based graph model employs a GAT as the foundational features encoder to facilitate cross-graph modal learning across different graph modalities. Given the two input features node representations $I^{\prime}=[I^{\prime}_{1},I^{\prime}_{2},...,I^{\prime}_{N}]\in\mathbb{R}^{N\times d}$ and $C=[C_{1},C_{2},...,C_{N}]\in\mathbb{R}^{N\times F}$, we can formulate the multi-modality attention coefficients in a GNN as:
@@ III-C Graph Attention Encoder
III-D Inter-Modality Feature Enhancement and Scaling
@@
III-E Cross-Graph Modal Fusion
To integrate features from cross-graph modality representations at varying scales, we concatenate each image embedding $E^{I}$ with its corresponding representation $\mathcal{H}^{I}$ to form $\hat{Z}^{I}=[\mathcal{H}^{I}\mathbin{\|}E^{I}]$, and similarly for clinical data, $\hat{Z}^{C}=[\mathcal{H}^{C}\mathbin{\|}E^{C}]$, resulting in final embedding matrices. We then combine these cross-modal embeddings through element-wise addition and apply a sigmoid function to create a common space: $\hat{Z}=\sigma(\hat{Z}^{I}+\hat{Z}^{C})$. This process aligns representations from both modalities in a common feature space. Furthermore, to integrate the feature spaces of images and meta-features in the same embedded space, we construct a similarity matrix $\textbf{S}\in\mathbb{R}^{N\times N}$ for each pair of similar patients using the final embedding $\hat{Z}$ learned from the model. The similarity between the $i$-th and $j$-th patients can be defined as follows:
@@ III-E Cross-Graph Modal Fusion
III-F Contrastive Loss Optimization
@@ IV Dataset Collection
V EXPERIMENTS
V-A Baseline Methods Comparison
@@
V-B Compare with the SOTA Methods
Compared with state-of-the-art (SOTA) methods, we examined the performance of the MHCA, DeAF, and TriFormer models, which employ 3D tensor-based CNNs as feature extractors combined with meta-features for multimodal training. Our experimental results yielded the following findings:
@@ VI Ablation Study
VI-A Parameters of $\beta$ in the Objective Function
@@
VI-B Multimodal Features Alignment in KL Divergence Converge
@@ VI-D Objective Function Ablation
VII Visualization of Multimodal Representation
VII-A Identification of PD subtype trajectories
@@
VII-B Multi-Class Trajectories in Melanoma Embedding
VIII Discussion
VIII-A Clustering Results
@@
VIII-B Neurodegenerative Subtype Analysis in PD
@@
VIII-C Explainable Modality Diagnosis